Sawyer Company had the following information for the year:

Direct materials used $ 194,100
Direct labor incurred (7,400 hours) $ 246,500
Actual manufacturing overhead incurred $ 315,000


Sawyer Company used a predetermined overhead rate using estimated overhead of $352,600 and 8,200 estimated direct labor hours. Assume the only inventory balance is an ending Finished Goods Inventory balance of $9,800. What was adjusted cost of goods sold?

  • $749,000

  • $745,800

  • $755,600

  • $758,800
Answer #1

Answer) $745800

Explanations:

1)Predetermined Overhead rate = 352600 / 8200 = $43 per DL

2) Overhead applied to Production = $43 x 7400 = $318200

3) Overapplied OH = $318200 - $315000 = $3200

4) Cost of good sold = Direct materials used + Direct labor incurred + Manufacturing overhead applied - Ending FG Inventory

= 194100 + 246500 + 318200 - 9800

= 749000

5) Adjusted Cost of goods sold = 749000 - Overapplied OH

= 749000 - 3200

= 745800